# Automorphisms of Plane Curves defined from Chebychev polynomials

### v1

@@ 3. Automorphism groups
Write $q:=p^{r}$. By the proof of Lemma 3.16, the condition $14d=q+$ 1 implies that $\varphi_{d}(x)$ splits completely in ${\mathbb{F}}_{q}[x]$. Moreover, the elements $x\mapsto-x$ and $x\mapsto(2x+12)/(2-x)$ (see Proposition 3.15) generate a subgroup $D_{3}\leq G/N$. As a consequence, Theorem 3.7 and Lemma 3.13 and Remark 3.14 imply that the quotient group $G/N\subset PGL(2,q)$ must be a dihedral group $D_{n}$ where $n$ divides $q\pm 1$ and $3|n$. Moreover, by Lemmas 3.8 and 3.12, one deduces that $n$ is odd.
@@
Assume $d>4$ and both $12d-$ 1 and $14d-$ 1 are not powers of the odd prime $p$. The curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ in characteristic $p$ has automorphism group isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z}$.
@@ 4. Nonisomorphic Maximal curves
Given a smooth, projective, and absolutely irreducible curve $X$ of genus $g$ over a finite field $\mathbb{F}_{q}$, the number of rational points satisfies the Hasse–Weil bound:
|#$|\#X(\mathbb{F}_{q})-(q+1)|\leq 2g\sqrt{q}.$
@@
From [3, Theorem 4.4] it is known that the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ is maximal over ${\mathbb{F}}_{q^{2}}$ if and only if $d$ divides $(q+1)/2$. In the special case $12d=q+$ 1, then Proposition 3.1 shows that the curve ${\mathcal{C}}_{d}$ is isomorphic to the Fermat curve ${\mathcal{F}}_{d}\colon x^{d}+y^{d}+z^{d}=0$ of degree $d=(q+1)/2$. Using Proposition 2.1, we know that if $d\neq(q+1)/2$, then the curve ${\mathcal{C}}_{d}$ has only $d$ total inflection points. This leads to a slightly different proof (but using the same ideas) of the following result presented in [1, Corollary 6.4]:
@@
Let $q$ be a power of a prime $p>3$ such that $\gcd(3,q+1)=1$. If $14n=q+$ 1, then for any divisor $m$ of $n$, the maximal curves ${\mathcal{C}}$ and $\mathcal{D}$ over $\mathbb{F}_{q^{2}}$ defined by

### v2

@@ 3. Automorphism groups
Write $q:=p^{r}$. By the proof of Lemma 3.16, the condition $4d=q+1$ implies that $\varphi_{d}(x)$ splits completely in ${\mathbb{F}}_{q}[x]$. Moreover, the elements $x\mapsto-x$ and $x\mapsto(2x+12)/(2-x)$ (see Proposition 3.15) generate a subgroup $D_{3}\leq G/N$. As a consequence, Theorem 3.7 and Lemma 3.13 and Remark 3.14 imply that the quotient group $G/N\subset PGL(2,q)$ must be a dihedral group $D_{n}$ where $n$ divides $q\pm 1$ and $3|n$. Moreover, by Lemmas 3.8 and 3.12, one deduces that $n$ is odd.
@@
Assume $d>4$ and both 2d12d-1 and 4d14d-1 are not powers of the odd prime $p$. The curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ in characteristic $p$ has automorphism group isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z}$.
@@ 4. Nonisomorphic Maximal curves
Given a smooth, projective, and absolutely irreducible curve $X$ of genus $g$ over a finite field $\mathbb{F}_{q}$, the number of rational points satisfies the Hasse Weil bound:
$|\#X(\mathbb{F}_{q})-(q+1)|\leq 2g\sqrt{q}.$
@@
From [3, Theorem 4.4] it is known that the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ is maximal over ${\mathbb{F}}_{q^{2}}$ if and only if $d$ divides $(q+1)/2$. In the special case $2d=q+1$, then Proposition 3.1 shows that the curve ${\mathcal{C}}_{d}$ is isomorphic to the Fermat curve ${\mathcal{F}}_{d}\colon x^{d}+y^{d}+z^{d}=0$ of degree $d=(q+1)/2$. Using Proposition 2.1, we know that if $d\neq(q+1)/2$, then the curve ${\mathcal{C}}_{d}$ has only $d$ total inflection points. This leads to a slightly different proof (but using the same ideas) of the following result presented in [1, Corollary 6.4]:
@@
Let $q$ be a power of a prime $p>3$ such that $\gcd(3,q+1)=1$. If $4n=q+1$, then for any divisor $m$ of $n$, the maximal curves $\mathcal{C}$} and $\mathcal{D}$ over $\mathbb{F}_{q^{2}}$ defined by
@@
Acknowledgment. The first author was partially supported by FAPESP grant No. 2023/08271-5, FAEPEX grant No. 2111/26 and CNPq grant No. 302774/2025 4.